OpenStudy (anonymous):

SIMPLIFY x/x^2 + 6x/ x-x^2 I need help..

OpenStudy (anonymous):

@tysonw @ganeshie8 @jim_thompson5910

OpenStudy (anonymous):

\[\frac{ x }{ x^2 }+\frac{ 6x }{ x(1-x) }\] Once you factor you can see the the common denominator is x²(1-x). Multiply the first fraction by (1-x)/(1-x) and the second by x/x to get a single denominator. \[\frac{ x(1-x) }{ x^2(1-x) }+\frac{ 6x(x) }{ x(1-x)(x) }\]

OpenStudy (anonymous):

\[\frac{ x-x^2+6x^2 }{ x^2(1-x)}\]

OpenStudy (anonymous):

can you simplify from here?

OpenStudy (anonymous):

yes give me sec :)

OpenStudy (anonymous):

-5x+6/(x-1) (x+1)

OpenStudy (anonymous):

not sure what you did. The denominator was already factored. Combine like terms in the numerator, then factor it.

OpenStudy (anonymous):

Can you step by step @peachpl

OpenStudy (anonymous):

Picking up from the last equation I posted